OpenStudy (anonymous):

4/3x +3/(3x+1) +2 =0 solve the equation algebraically/show steps.

OpenStudy (lalaly):

\[\frac{4}{3x}+\frac{3}{(3x+1)}+2=0\]

OpenStudy (lalaly):

is this the equation?

OpenStudy (anonymous):

yes :)

OpenStudy (lalaly):

first multiple both sides by 3x\[3x(\frac{4}{3x} + \frac{3}{(3x+1)}+2)= 0 \times 3x\]\[4+\frac{9x}{(3x+1)} + 6x =0\]Now multiply both sides by (3x+1) \[4(3x+1)+9x+6x(3x+1)=0\]Expand the brackets\[12x+4+9x+18x^2+6x=0\]

OpenStudy (lalaly):

\[18x^2+27x +4 =0\]

OpenStudy (lalaly):

(3x+4)(6x+1) =0 x=-4/3 x= -1/6

OpenStudy (anonymous):

ah ok, looks like i was just messing up the factoring. Thanks a bunch for showing the steps for me--I couldn't figure out which part I'd messed up on. :)

OpenStudy (lalaly):

your welcome =D

OpenStudy (anonymous):

Have a good one! :D